# What is glacier sliding?

### v1

@@ 3.7 Frozen-thawed transitions and sub-temperate sliding
In some simpler theories both sliding and slip are considered to be negligible for frozen beds (Lliboutry,, 1966; Cuffey and Paterson,, 2010, 7.1.1), but most models either simply neglect a temperature control on sliding (including all models inverting for basal traction with present-day velocity fields in ISMIP6 experiments) or take a heuristic approach to scaling basal traction with temperature to avoid a stress singularity at the frozen thawed boundary (e.g. Tarasov and Peltier,, 2007; Mantelli and others,, 2019) giving
@@
as an adaptation of Eq. 7 where $T_{h}=T_{b}-T_{pmp}$ is the homologous temperature calculated from the pressure melting point, $T_{pmp}$, and ice temperature at the base, $T_{b}$. This corroborates with field studies finding slow (<0.2 m a-1) slip rates even with temperatures some degrees below freezing for both soft- and hard-bed settings (Hallet and others,, 1986; Echelmeyer and Zhongxiang,, 1987; Cuffey and others,, 1999; Fitzsimons and others,, 1999). The mechanical process or processes responsible for sub-temperate sliding in these instances remains uncertain (Cuffey and Paterson,, 2010, 7.2.8) but may be related to a very thin (nm) premelt layer at the ice bed (Shreve,, 1984), the inclusion of solid friction (Fowler,, 1986), or sediment deformation despite freezing conditions Waller, (2001). Separately, form drag will still be expected under freezing bed conditions, though at a slower rate than under equivalent stress conditions in colder ice. This processes is framed by Weertman, (1967) as the continuation of enhanced creep despite the cessation of regelation.Expanding sub-temperate sliding out to a grid cell under consideration also allows for the situation of both freezing and thawed areas within the same grid cell (Fig. 7), which may be a common situation in frozen-thawed transition zones (e.g. Oswald and Gogineni,, 2012; Dawson and others,, 2024). Generally, a pattern of greater basal traction in interior ice-sheet regions where a frozen bed is considered more likely is obtained in ice-sheet wide inversions (Larour and others,, 2012; Morlighem and others,, 2013).
@@ 4 Subglacial Hydrology
Subglacial hydrology and basal sliding are intimately linked in many glacial settings (see reviews by Irvine-Fynn and others,, 2011; Ashmore and Bingham,, 2014, and Nienow and others,, 2017). Importantly, water pressure influences sliding velocity via the effective pressure $N$ and sliding velocity can in turn influences $N$ as changes in cavitation influence the hydraulic conductivity and connectivity of the drainage system (e.g. Schoof,, 2010). The subglacial hydrological system is also important for transporting and evacuating glacial melt water, sediment, and nutrients (Swift and others,, 2005; Death and others,, 2014; Hawkings and others,, 2015; Hofstede and others,, 2023) – each of which can be a product of sliding, via frictional heat dissipation or subglacial erosion.
@@
In most numerical models, subglacial hydrology is either absent or is parameterised in a basic fashion through the effective pressure $N$ term in the sliding relationship (Table. 1, Fig. 9) – with at least three materially different ways of including $N$ in Eq. 1 (Table 1). While in the physical sense the meaning of $N$ as subglacial water pressure subtracted from the overburden pressure at a point in time and space is very clear, neither its meaning nor ways of quantifying a representative $N$ at a given scale are straight forward in a modelling sense. Subglacial hydrological environments are very spatially diverse even at scales of tens of metres, with various terms (e.g. ‘channelised; distributed; partly-connected’) referring to various end members in a continuum of characteristics (e.g. Hubbard and Nienow,, 1997; Hoffman and others,, 2016; Rada and Schoof,, 2018). Borehole observations evidence this complexity, with widely varying $N$ time series observed in adjacent boreholes (Doyle and others,, 2018; Rada and Schoof,, 2018; Doyle and others,, 2021) capturing distinctly separate components of subglacial hydrology, consistent with theoretical studies finding low-pressure conduits adjacent to higher-pressure regions with reduced hydrological connectivity (Engelhardt and Kamb,, 1997; Hewitt and others,, 2018). Moreover, subglacial drainage systems are observed to evolve diurnally and seasonally in response to supraglacial melt water inputs (Hubbard and Nienow,, 1997), even beneath ice $\sim$1 km thick (Chandler and others,, 2013; Doyle and others,, 2021). This evolution in turn leads to diurnal and seasonal variability in sliding (Willis,, 1995; Zwally and others,, 2002; Bartholomew and others,, 2011; Nienow and others,, 2017). Such temporal variability is not captured by ice sheet or glacier models employing temporally constant parameters in the friction law.
While the influence of $N$ is fairly well understood for specific sliding sub-processes in isolation – for example, cavitation over a relatively flat bedrock area (Helanow and others,, 2021), and sediment shearing in a laboratory setting (e.g. Iverson and others,, 1998) or through a numerical continuum model (e.g. Damsgaard and others,, 2020) – the influence of $N$ over a large area encompassing a diverse set of sliding sub-processes is not well constrained. Additionally, it is not comprehensively established how $N$ should be incorporated within Eq. LABEL:eq:1, whether through a linear multiplication (e.g. Zoet and Iverson,, 2020; Choi and others,, 2022), raised to a power (Budd and others,, 1984; Budd and Jacka,, 1989), or included in a more complicated manner (Helanow and others,, 2021) – a situation that may be more uncertain if the influence of rough topography is also incorporated.
@@
Despite the importance of subglacial hydrology as a control on sliding at sub-annual scales, its overall influence on long-term glacier and ice sheet evolution remains debated, particularly with respect to future changes of the Greenland Ice Sheet (e.g. Tedstone and others,, 2015; Doyle and others,, 2015; Beckmann and Winkelmann,, 2023), and Antarctic ice streams (Bougamont and others,, 2015), with major changes in glacier and ice-sheet velocity fields typically linked to geometry changes and ocean forcing rather than intrinsic changes in the hydrological system whether or not substantial surface melt occurs (e.g. Vincent and others,, 2012; Habermann and others,, 2013; Moon and others,, 2014; Cook and others,, 2016; Seroussi and others,, 2017; Catania and others,, 2018; Dehecq and others,, 2018; Gimbert and others,, 2021; Frank and others,, 2022) with the exception of surging glaciers (Sevestre and Benn,, 2015). However, large-scale observations do indicate that major changes in the hydrological system – such as a switch from marine-terminating to land-terminating outlet glaciers, or an advancement of the ablation zone – may influence large-scale geometry and velocity patterns over a multi-annual timescale (Maier and others,, 2022). It may therefore be reasonable to suggest that diurnal and seasonal hydrology variations are of secondary importance to glacier geometry and ocean forcing at the annual and greater timescales important for future predictions. The exception to this is that the subglacial water pressure of marine-terminating outlet glaciers grounded below sea level is likely significantly controlled by their proximity to the ocean, with inferred very low basal traction values and very high subglacial water pressure very close to overburden (e.g. Stearns and Van Der Veen,, 2018; Minchew and others,, 2019).
@@ 5 Heuristic studies
A separate approach to investigating glacier sliding is to use large-scale remotely sensed or field data to heuristically (i.e. through process-agnostic iterative experimentation and adjustment) determine the form of the sliding relationship, which may then be compared to theoretical expectations (Section 3). A time series with significant variation in surface geometry is best suited to the decadal time scales over which mass loss is most often considered (Gillet-Chaulet and others,, 2016; Gilbert and others,, 2023), but several other variations have been exploited (Table 2). Considered in their entirety, these studies indicate that variation in the form of the sliding relationship across variable glacier settings cannot be captured by a single tunable coefficient, $C$, obtained through a standard inversion procedure. For example, Gimbert and others, (2021) and Gilbert and others, (2023) find $m=3.1\pm 0.3$ in Weertman-type sliding for Argentière Glacier in the French Alps using a combination of surface velocity changes, length changes, and direct sliding measurements while Gillet-Chaulet and others, (2016) find $m\leq 5$ and up to 50 well reproduces surface velocities at Pine Island Glacier, Antarctica. Maier and others, (2021) meanwhile indicate that a single $m$ value, also for Weertman-type sliding, cannot be applied to the entire Greenland Ice Sheet with, for example, an $m$ of 8 in northeast Greenland contrasting an $m$ of 4 in central southwest Greenland. Note that in standard applications using Weertman-type sliding, $C$ is the subject of the inversion and varies in space, while $m$ is kept constant.
@@
Such studies offer a testing ground for proving the efficacy of a given sliding relationship, but it is important to note that a match to a given theory does not on its own demonstrate that the physical processes that build it up are actually occurring in the study location. This is straightforward in the case of $m\leq 5$ from Gillet-Chaulet and others, (2016) at Pine Island Glacier. No existing theory postulates this relationship (Section 3) so the utility of $m\leq 5$ in Weertman-type sliding comes either from the fact that it emulates regularised-Coulomb sliding that does have a large body of theory behind it, or it represents hitherto unappreciated processes not presently incorporated into a process-based sliding relationship. Meanwhile, it is difficult to interpret the findings of Minchew and others, (2016) from a setting known to have significant till coverage Björnsson and others, (2003) as anything other than plastic deformation of till. In contrast, Gimbert and others, (2021) and Gilbert and others, (2023) suggest that their findings of $m\sim 3$ reflects only the ice deformation component of the original Weertman, (1957) paper without the regelation component where $m=1$, neglecting the fact that the original Weertman theory is somewhat dependent upon a combination of both regelation and ice deformation, and making no allowance for the slip component actively observed beneath Argentière (Vincent and Moreau,, 2016; Gimbert and others,, 2021). Maier and others, (2021) also invoke Weertman-type hard-bed physics as justification for lower values of $m$, despite the fundamental issues underlying the theory (Section 3.2) and the fact that no theory building upon Weertman, (1957), or the original study, provide a physical justification for $n=4$ or above.
Taken collectively, heuristic studies emphasise a heterogeneity between settings that is not easily reconciled with a single sliding relationship controlled by a single tunable parameter. However, they also present perhaps the most definitive method of accurately determining the optimum sliding relationship for a given modelling setting. We suggest that development of a standardised methodology that can be efficiently applied to varying settings utilizing increasingly long datasets will be of great use in transparently tailoring sliding relationships for given modelling applications.
@@ 6 Discussion
6.1 General differences between sliding in glaciers and ice-sheets
By convention, the sliding of ice sheets is treated in the same manner as the sliding of valley glaciers from which foundational sliding theories are typically derived (e.g. Hodge,, 1974; Engelhardt and others,, 1978; Fischer and Clarke,, 1997; Willis,, 1995; Bouchayer and others,, 2024). However, differences intrinsic to the two settings, and choices made in model development, may significantly influence the combination of sliding sub-processes occurring and those represented within a given grid cell. Valley glaciers (i.e. those occupying a limited geographical area, not draining a large accumulation area ice field, and comprising only a handful of tributaries) have several features distinguishing them from both ice-sheet interiors and ice-sheet outlet glaciers. First, valley glaciers generally occupy a sufficiently small area such that, as a rule, major geological boundaries are not crossed beneath glacier, making a distinction between soft- and hard-bed sliding more reasonable (and indeed consequential, with surging glaciers generally occurring over inferred soft beds; Cuffey and Paterson,, 2010, Section 12.2). Second, valley glaciers tend to be much steeper – though the distribution of driving stresses is similar (Meyer and others,, 2018) – with underlying topography in the mountainous regions where valley glaciers are present typically rough, compared to the more mixed geomorphology that underlies present and palaeo ice sheets. Third, valley glaciers generally do not reach the high velocities that characterize the large outlet glaciers important for dynamic changes in ice sheets with velocities only exceeding 100 m a-1 for large glaciers comprising an extensive accumulation area and with several tributaries (Laumann and Wold,, 1992; Jouvet and Huss,, 2019). Fourth, if valley glaciers do terminate in lakes or the ocean then the glacier immediately inland tends to be much steeper than ice-sheet outlet glacier counterparts meaning the hydrological influence of the water body also decreases rapidly inland.
The net effect of these factors, and the generally much greater resolution in modern models of valley glaciers (cf. Aschwanden and others,, 2019; Gilbert and others,, 2020) is that that the processes represented by a single grid area, $R$, can differ substantially between glaciers and ice sheets (Fig. 2). This is born out in the limited evidence from heuristic studies (Section 5) where all valley glaciers investigated in this manner are found to be better represented by unbounded sliding relationships (Table 2). In fact, we are not aware of valley glaciers being represented by a regularised-Coulomb relationship in any instance. Care is needed therefore, in applying the findings of sliding studies conducted at valley glaciers (Gimbert and others,, 2021; Gilbert and others,, 2023) to ice sheets, and vice versa (Tsai and others,, 2021), in the same manner that care is needed in applying laboratory and numerical modelling studies to glacier and ice sheet settings, even though the individual sub-processes will remain valid in both settings.

### v3

@@ 3.7 Frozen-thawed transitions and sub-temperate sliding
In some simpler theories both sliding and slip are considered to be negligible for frozen beds (Lliboutry,, 1966; Cuffey and Paterson,, 2010, 7.1.1), but most models either simply neglect a temperature control on sliding (including all models inverting for basal traction with present-day velocity fields in ISMIP6 experiments) or take a heuristic approach to scaling basal traction with temperature to avoid a stress singularity at the frozen thawed boundary (e.g. Fowler and Larson,, 1978; Tarasov and Peltier,, 2007; Mantelli and others,, 2019) giving
@@
as an adaptation of Eq. 1 where $T_{h}=T_{b}-T_{pmp}$ is the homologous temperature calculated from the pressure melting point, $T_{pmp}$, and ice temperature at the base, $T_{b}$. This corroborates with field studies finding slow (<0.2 m a-1) slip rates even with temperatures some degrees below freezing for both soft- and hard-bed settings (Hallet and others,, 1986; Echelmeyer and Zhongxiang,, 1987; Cuffey and others,, 1999; Fitzsimons and others,, 1999). The mechanical process or processes responsible for sub-temperate sliding in these instances remains uncertain (Cuffey and Paterson,, 2010, 7.2.8) but may be related to a very thin (nm) premelt layer at the ice bed (Shreve,, 1984), the inclusion of solid friction (Fowler,, 1986; Section 3.2), or sediment deformation despite freezing conditions Waller, (2001). Separately, both form drag and geometric resistance will still be expected under freezing bed conditions, though at a slower rate than under equivalent stress conditions in colder ice. This processes is framed by Weertman, (1967) as the continuation of enhanced creep despite the cessation of regelation, and matches Eq. 8 with an absent or much reduced $\boldsymbol{\mathcal{S}}_{T}$ component. Expanding sub-temperate sliding out to a grid cell under consideration also allows for the situation of both freezing and thawed areas within the same grid cell (Fig. 8), which may be a common situation in frozen-thawed transition zones (e.g. Oswald and Gogineni,, 2012; Dawson and others,, 2024). Generally, a pattern of greater basal traction in interior ice-sheet regions where a frozen bed is considered more likely is obtained in ice-sheet wide inversions (Larour and others,, 2012; Morlighem and others,, 2013).
@@ 4 Subglacial Hydrology
Subglacial hydrology and basal sliding are intimately linked via close feedbacks in many glacial settings, with water pressure influencing sliding velocity via the effective pressure $N$ and sliding velocity in turn influencing $N$ as changes in cavitation affect hydraulic conductivity and drainage system connectivity (e.g. Schoof,, 2010). There are too many facets of this topic to cover comprehensively here (reviews include: Irvine-Fynn and others,, 2011; Ashmore and Bingham,, 2014; Chu,, 2014; Siegert and others,, 2016, Nienow and others,, 2017; Davison and others,, 2019), so we focus briefly on the importance of a spatio-temporally varying subglacial hydrology system, without considering the mechanisms themselves in detail.
@@
In most numerical ice sheet and glacier models, subglacial hydrology is either absent or is parametrised in a simple fashion through the effective pressure $N$ term in the sliding relationship (Table. 1, Fig. 9) – with at least three materially different ways of including $N$ in Eq. 1 (Table 1). While in the physical sense the meaning of $N$ as subglacial water pressure subtracted from the overburden pressure at a point in time and space is very clear, its meaning, and ways of quantifying a representative $N$ at a given scale are not straight forward in a modelling sense. Subglacial hydrological environments are very spatially diverse even at scales of tens of metres, with various terms (e.g. ‘channelised; distributed; partly-connected’) referring to various end members in a continuum of characteristics (e.g. Hubbard and Nienow,, 1997; Hoffman and others,, 2016; Rada and Schoof,, 2018). Borehole observations and investigations of deglaciated terrains (Fig. 9) evidence this complexity, with widely varying $p_{w}$ time series observed in adjacent boreholes (Doyle and others,, 2018; Rada and Schoof,, 2018; Doyle and others,, 2021) capturing distinctly separate components of subglacial hydrology, consistent with theoretical studies finding low-pressure conduits adjacent to higher-pressure regions with reduced hydrological connectivity (Engelhardt and Kamb,, 1997; Hewitt and others,, 2018). Therefore, while the influence of $N$ is fairly well understood for specific sliding sub-processes in isolation – for example, cavitation over a relatively flat bedrock area (Helanow and others,, 2021), and sediment shearing in a laboratory setting (e.g. Iverson and others,, 1998) or through a numerical continuum model (e.g. Damsgaard and others,, 2020) – it remains to fully quantify the influence of a spatially variable $N$ within a given model grid cell. We are so far only aware of the study of Tsai and others, (2021) which locally sets $\tau_{b}$ to zero over an area considered to be comprised of cavities, or elsewise calculates it following a power-law relationship, resulting in a relationship that overall shares similarities with the Budd and Jacka, (1989) sliding relationship (Table 1).
Beyond this, it is challenging to know exactly how to relate $N$ to other parameters and variables within a standard sliding relationship. In our framework (Section 2), varying $N$ could be incorporated within the parameter $P_{R_{i}}$ in Eq. 9 or power-law equivalent, though calculating its actual spatial distribution would still remain a major challenge. Most extant cases implement a simple linear multiplication (e.g. Choi and others,, 2022), but what, for example, if the traction within a region with low $N$ for a large proportion of its area was actually controlled by a small prominent pinning-point with high $N$?
@@
Moreover, subglacial drainage systems are also observed to evolve diurnally and seasonally in response to supraglacial melt water inputs (Hubbard and Nienow,, 1997), even beneath ice ca. 1 km thick (Chandler and others,, 2013). This evolution in turn leads to diurnal and seasonal variability in sliding (Willis,, 1995; Zwally and others,, 2002; Bartholomew and others,, 2011; Nienow and others,, 2017). Such temporal variability in $N$ is by design not captured by ice sheet or glacier models employing temporally constant parameters for the basal traction relationship. However, despite the clear importance of subglacial hydrology as a control on sliding at sub-annual scales however, and with the exception of surging glaciers (Sevestre and Benn,, 2015), its overall influence on long-term glacier and ice sheet evolution remains debated, particularly with respect to future changes of the Greenland Ice Sheet (e.g. Tedstone and others,, 2015; Doyle and others,, 2015; Beckmann and Winkelmann,, 2023), with major and sustained changes in glacier and ice-sheet velocity fields typically linked to geometry changes rather than intrinsic changes in the hydrological system (e.g. Vincent and others,, 2012; Habermann and others,, 2013; Moon and others,, 2014; Cook and others,, 2016; Seroussi and others,, 2017; Catania and others,, 2018; Dehecq and others,, 2018; Gimbert and others,, 2021; Frank and others,, 2022). However, large-scale observations do indicate that major changes in the hydrological system – such as a switch from marine-terminating to land-terminating outlet glaciers, or an advancement of the ablation zone – may influence large-scale geometry and velocity patterns over a multi-annual timescale (Maier and others,, 2022). Subglacial water pressure also plays a crucial control at marine-terminating outlet glaciers grounded below sea level, which we do not cover in detail here, where very low inferred basal traction values and subglacial water pressure very close to overburden make systems particularly susceptible to changes in $N$ (e.g. Katz and Worster,, 2010; Stearns and Van Der Veen,, 2018; Minchew and others,, 2019; Bradley and Hewitt,, 2024).
@@ 5 Heuristic studies
A separate approach to investigating glacier sliding is to use large-scale remotely sensed or field data to heuristically (i.e. through process-agnostic iterative experimentation and adjustment) determine the form of the sliding relationship, which may then be compared to theoretical expectations (Section 3, Table 1). A time series with significant variation in surface geometry is best suited to the decadal time scales over which mass loss is most often considered (Gillet-Chaulet and others,, 2016; Gilbert and others,, 2023), but several other variations have been exploited (Table 2). Considered in their entirety, these studies indicate that variation in the form of the sliding relationship across variable glacier settings is not readily captured by a single tunable coefficient, $C$, obtained through a standard inversion procedure. For example, Gimbert and others, (2021) and Gilbert and others, (2023) find $m=3.1\pm 0.3$ in power-law sliding for Argentière Glacier in the French Alps using a combination of surface velocity changes, length changes, and direct sliding measurements while Gillet-Chaulet and others, (2016) find $m\leq 5$ and up to 50 well reproduces surface velocities at Pine Island Glacier, Antarctica. Maier and others, (2021) meanwhile indicate that a single $m$ value, also for power-law sliding, cannot be applied to the entire Greenland Ice Sheet with, for example, an $m$ of 8 in northeast Greenland contrasting an $m$ of 4 in central southwest Greenland. Note that in standard applications using power-law sliding, $C$ is the subject of the inversion and varies in space, while $m$ is kept constant.
@@
West Greenland – outlet glacier Front position from 1985-2019 Jager and others, (2024) Regularised-Coulomb
@@
Such studies offer a testing ground for proving the efficacy of a given sliding relationship, but it is important to note that a match to a given theory does not on its own demonstrate that the physical processes that build it up are actually occurring in the study location. This is straightforward in the case of $m\leq 5$ from Gillet-Chaulet and others, (2016) at Pine Island Glacier. No existing theory postulates this relationship (Section 3) so the utility of $m\leq 5$ in power-law sliding comes either from the fact that it emulates regularised-Coulomb sliding that does have a large body of theory behind it, or it represents a hitherto unappreciated set of processes not presently incorporated into a process-based sliding relationship. Meanwhile, it is difficult to interpret the findings of Minchew and others, (2016) from a setting known to have significant till coverage (Björnsson and others,, 2003) as anything other than plastic deformation of till. In contrast, Gimbert and others, (2021) and Gilbert and others, (2023) suggest that their findings of $m\sim 3$ reflects only the ice deformation component of the original Weertman, (1957) paper without the regelation component where $m=1$, neglecting the fact that the original Weertman theory is somewhat dependent upon a combination of both regelation and ice deformation, and not making an allowance for the slip component actively observed beneath Argentière (Vincent and Moreau,, 2016; Gimbert and others,, 2021). Maier and others, (2021) also invoke Weertman-type hard-bed physics as justification for lower values of $m$, despite the fundamental issues underlying the theory (Section 3.2) and the fact that no theory building upon Weertman, (1957), or the original study, presently provide a physical justification for $n=4$ or above.
Taken collectively, heuristic studies emphasise a heterogeneity between settings that is not easily reconciled with a single sliding relationship controlled by a single tunable parameter. However, they also present perhaps the most definitive method of accurately determining the optimum sliding relationship for a given modelling setting. A standardised methodology that can be efficiently applied to varying settings utilizing increasingly long datasets could therefore be of use in transparently tailoring sliding relationships for given modelling applications.
@@ 6 Discussion
6.1 Applying an inner-outer flow framework to multiple sub-processes: bridging hard and soft beds and highlighting the importance of topographic roughness
We can separate the sub-processes covered in Section 3 that deal with continuous, rather than stick-slip, motion into the three categories defined for Eq. 8. All soft-bed processes and hard-bed processes over macroscopically flat surfaces (Sections 3.1, 3.2) can be defined as slip processes (within $\boldsymbol{\mathcal{S}}_{T}$). All processes concerning viscous deformation of ice (be it cold, the basal ice layer, or temperate, Sections 3.3, 3.4) can be defined as viscous deformation processes (within $\boldsymbol{\mathcal{D}}_{T}$). Hard-bed sliding processes involving ice deformation straddle slip, deformation, and geometric resistance. The most precise way to include these would be to deconstruct their components (i.e. free slip at the ice-bed interface and over cavities, geometric resistance from obstacles, enhanced deformation) and reimpose them as segregated sub-processes, but this would then place the onus of calculating their evolution with increasing $u_{b}$ on each usage rather than allowing for it to be more simply sub-parametrised. This creates particular complications for cavitation, which is very computationally expensive to simulate (Helanow and others,, 2021). For the purposes of this discussion we therefore treat ostensibly ‘hard-bedded’ power-law and regularised-Coulomb theories (Section 3.2) as slip processes within $\boldsymbol{\mathcal{S}}_{T}$. This involves the explicit assumption that the region $R$ is then considered as sufficiently large so as to accommodate topographic features with wavelengths >25 m, if present.
The relative balance of $\boldsymbol{\mathcal{S}}_{T}$, $\boldsymbol{\mathcal{D}}_{T}$, and $\boldsymbol{\mathcal{G}}$ within Eq. 8 then allows us to make assessments of their influence on the overall traction at the inner-outer boundary for a given setting at the scale of ice-sheet or glacier model discretisation. If, as appears reasonable for relatively smooth isolated bed sub-sections (Section 3.6), regularised-Coulomb can represent slip over both soft and hard beds in the presence of abundant meltwater, the term $\boldsymbol{\mathcal{S}}_{T}$ becomes agnostic over whether the bed is soft or hard. Then, if the region is relatively planar and $b_{io}-b_{b}$ is relatively small, $\boldsymbol{\mathcal{S}}_{T}$ dominates and it is most likely that a regularised-Coulomb relationship emerges. However, if $\boldsymbol{\mathcal{D}}_{T}$ (determined as a power-law, Eq. 10) and $\boldsymbol{\mathcal{S}}_{T}$ (related to increasing velocity, obstacle prominence, and obstacle frequency, Eq. 13) dominate, then the traction for the given region will be better represented through a power-law. This implies that bed roughness, not bed type is the crucial control on the form of the sliding relationship.
The case of frozen-bed sliding would furthermore shut off, or greatly reduce the importance of $\boldsymbol{\mathcal{S}}_{T}$, resulting in power-law sliding regardless of the roughness characteristics of the region in consideration.
6.2 Maximum cavitation sizes, and revisiting Iken’s bound
[…]
If large topographic obstacles are present, such as those in Fig. 10, but large-scale cavities are not concomitant (Fig. A3), then the given grid cell area may come to be dominated by topographic roughness rather than having cavitation essentially smooth over the ice-bed interface (Fig. A3). As discussed in Section 6.1, the stage is then set for the traction defined at the inner-outer boundary to be dominated by resistance related to viscous deformation ($\boldsymbol{\mathcal{D}}$) and bed geometry ($\boldsymbol{\mathcal{G}}$), rather than a bounded basal traction relationship overall. Nonetheless, while we can place common-sense limits on cavity size in deglaciated areas, a full understanding of the scale distribution of cavities, and its bearing upon Iken’s bound, is an important future research goal.
In addition, the term $\mathrm{tan}(\beta)$ in Iken’s bound (Eq. 14) will tend towards infinity as slopes opposing the prevailing approach perpendicular , and it is therefore no longer a bound in the situation of steeply oriented cliff faces, as seen in Fig. 10 and in many rough deglaciated landscapes. We suggest that the unsuitability of Iken’s bound in regions featuring steep slopes, coupled with the importance of large topographic obstacles in increasing the $\boldsymbol{\mathcal{D}}_{T}$ and $\boldsymbol{\mathcal{G}}$ terms in Eq. (8) at larger scales may explain the continued efficacy of a power-law sliding relationships in ice-sheet and glacier settings (Table 2).
Figure 10: Rough topography near Hardangerfjord, western Norway. Hut location is 60.4526 N, 6.1406 E, inset shows ArcticDEM topography for the valley between the hut and the small glacier in the distance with no vertical exaggeration (Porter and others,, 2018). Palaeo flow direction at around 20 kyr ago available from e.g. Mangerud and others, (2019); Jungdal-Olesen and others, (2024). Numerous very steep surfaces opposing the direction of palaeo ice sheet motion can be seen at both a micro and macro scale. Photo taken by RL in late September 2024.
6.3 Does a unified basal sliding relationship exist?
A defining objective of much research into sliding is to uncover a ‘universal’ sliding relationship which can provide an accurate description of sliding processes across all settings, scales, and time periods. Generally, the implicit assumption is that this will be controlled by a single tunable coefficient (henceforth single coefficient) and/or calculated subglacial water pressure (e.g. Minchew and Joughin,, 2020b). Unfortunately, while it may be possible to describe glacier sliding by incorporating a multitude of stress contributions (Eq. 8) the complex nature of myriad sliding sub-processes and their scale and setting dependence (Section 3) likely precludes a single-coefficient relationship applicable to all settings – an outlook supported by the diversity of optimum single-coefficient relationships in Table 2. Instead, the studies reviewed here and Eq. 8 point towards a compound relationship that bears closer resemblance to descriptions of ice strain comprising several deformation mechanisms (e.g. Eq. 3 of Goldsby and Kohlstedt,, 2001).
Nonetheless, we may be able to talk about the general nature of such a compound relationship if, as in Section 6.1, we take regularised-Coulomb (Eq. 9) as the only constitutive relationship for $\boldsymbol{\mathcal{S}}_{T}$, and then take Eq. 10 as the only constitutive relationship for $\boldsymbol{\mathcal{D}}_{T}$. This allows us to take $\boldsymbol{\mathcal{S}}_{T}$ and $\boldsymbol{\mathcal{D}}_{T}$ as single terms, rather than sums as
[…]
and
[…]
where the subscript $\,{}_{Ta}$ refers to total aggregate $P_{R_{r}}$, and ${u_{t}}_{R_{r}}$, ${m_{R_{r}}}$, $n_{R_{r}}$, and $K_{R_{r}}$ are the representative parameters for the given region. $P_{R_{r}}$ and $K_{R_{r}}$ can be defined as e.g.
$P_{R_{r}}(x,y)=\frac{1}{A_{R}}\sum_{i\in\{\alpha,\beta,\gamma,...\}}I_{R_{i}}(x,y)P_{R_{i}}A_{R_{i}}\,$ (19)
for $P_{R_{r}}$ where $A_{R_{i}}$ is the area of each sub region. $m_{R_{i}}$ and $n_{R_{i}}$ cannot be calculated as a weighted average, reducing the set to a single power is a reasonable approximation in keeping with the use of $n=3$ over all glacier and ice sheet settings. Nonetheless, careful consideration should be given to $n_{R_{i}}$ in particular, given its potential dependence on whether ice in the inner layer is temperate, or has a high sediment content (Section 3.3, 3.4). Keeping the definition for $\boldsymbol{\mathcal{G}}$, but integrating over $R$ gives
[…]
where $\boldsymbol{\mathcal{G}}_{a}$ refers to the aggregate geometric resistance.
If we assume that each $\boldsymbol{\mathcal{G}}_{a}$, $\boldsymbol{\mathcal{S}}_{T_{a}}$, $\boldsymbol{\mathcal{D}}_{T_{a}}$ then defines a representative traction component at each coordinate following a moving window average, rather than for within a specific region, and with an $R$ size that at least approaches a representative surface element (Section 1), we can set
$\boldsymbol{\tau}_{b}=\boldsymbol{\mathcal{S}}_{Ta}+\boldsymbol{\mathcal{D}}_{Ta}+\boldsymbol{\mathcal{G}}_{a}$ (21)
which can be controlled by a total of 7 parameters (though dependent on the exact nature of the function used to describe $\boldsymbol{\mathcal{G}}_{a}$). Further, in the condition that $\boldsymbol{\mathcal{G}}_{a}$ does not dominate $\boldsymbol{\tau}_{b}$ in a complicated manner (or in fact follows a power-law relationship similar to $\boldsymbol{\mathcal{D}}_{T_{a}}$), and following the other assumptions in this section, we can say that $\boldsymbol{\tau}_{b}$ should always be defined as no more than the sum of a regularised-Coulomb and power-law relationship. For the one-dimensional case aligning with Table 1 that is
$\tau_{b}=f_{rC}(u_{b})+f_{pl}(u_{b})$ (22)
where $f_{rC}(u_{b})$ represents regularised-Coulomb sliding of the form Eq. 17 and $f_{pl}(u_{b})$ is a power-law of the form Eq. 18. Effective pressure may be implicitly included in Eqs. 21, 22 through parameters in $\boldsymbol{\mathcal{S}}_{Ta}$ and $f_{rC}(u_{b})$ but is presently missing as an explicit parameter here. From the perspective of the inner-outer flow boundary, the relative contribution of $\boldsymbol{u}_{b}$ to $\boldsymbol{u}_{s}$ and $\dot{\boldsymbol{\epsilon}}$ respectively will also be a function of landscape, ice rheology, and parameters for slip at the ice-bed interface. We note $f_{pl}(u_{b})$ will not be directly related to $N$ and should therefore remain important in areas with high $\dot{\boldsymbol{\epsilon}}$ within the inner flow at both low and high water pressures. Improved understanding of $\boldsymbol{\mathcal{G}}_{a}$ would allow improved confidence regarding the applicability of Eq. 22.
However, as standard inversion procedures are not well-suited to determining multiple parameters simultaneously (Morlighem and others,, 2010) and as a compound sliding relationship would complicate implementation, the choice between a compound friction law (e.g. Eq. 21) and single-term equation from Table 1 becomes a compromise between a more realistic representation of friction and a less practical numerical implementation. It may therefore be more appropriate to select a flexible yet simple equation as a sliding relationship which is capable of representing a wide range of behaviours.